OpenStudy (anonymous):

How do I set this up? Please help: Find the nth term of a sequence whose first several terms are given. 2, 11, 20, 29,...

OpenStudy (anonymous):

looks like the pattern is to add 9

OpenStudy (anonymous):

okay we know that is it an arithmetic sequence with a difference of +9 and a1=2 so c=2-9=-7 the formula is an=dn+c an=9n-7

OpenStudy (anonymous):

Thanks!